OpenStudy (anonymous):

what 2 factors of -28 equals -8

hartnn (hartnn):

why do u need them ?

OpenStudy (anonymous):

do not know the answer

OpenStudy (solomonzelman):

\[-28=~(-1) \times 2 \times 2 \times 7\]\[-8=(-1) \times 2 \times 2 \times 2\]

OpenStudy (solomonzelman):

you can brake it up into factors differently, you would have a couple of answers.

hartnn (hartnn):

i asked because there are no such integers